# Feedback-driven quantum reservoir computing for time-series analysis

### v1

@@ III.1 Short-term memory capacity
Figure 3: (a) Reservoir performance for the short-term memory task $R^{2}_{d}$ plotted as a function of the delay $d$ with $a_{\mathrm{fb}}=2.0$ (blue), $a_{\mathrm{fb}}=4.0$ (green), $a_{\mathrm{fb}}=6.0$ (purple), and $a_{\mathrm{fb}}=8.0$ (yellow). (b) The dependency of the total capacity $C_{\Sigma}$ on the feedback strength $a_{\mathrm{fb}}$. Both $R^{2}_{d}$ and $C_{\Sigma}$ are averaged over $128$ samples with respect to the $U_{\mathrm{res}}$.
Figure 3(a) exhibits the memory performance $R^{2}_{d}$ for varying feedback weights $a_{\mathrm{fb}}$, with the input weight fixed at $a_{\mathrm{in}}=0.001$. At a feedback strength of $a_{\mathrm{fb}}=2.0$, $R^{2}_{d}$ becomes close to $1$ for delays up to $d=4$, followed by a gradual decline for longer delays. Considering that the quantum state of the $k$-th cycle is determined only by $s_{k}$ and $\bm{z}_{k-1}$, this robust memory performance indicates that the input histories of $\{s_{k}\}$ are adequately accumulated within the measurement results $\bm{z}_{k-1}$. Essentially, the quantum reservoir effectively extracts key features of the provided dataset through the read-out measurements, and these features are succeeded over different cycles via the feedback connections. In this manner, our feedback-driven QRC overcomes the alternation of the quantum state caused by the measurement back-action, thereby acquiring the fading-memory property.
@@
Figure 3(b) depicts the total capacity $C_{\Sigma}$ as a function of the feedback strength $a_{\mathrm{fb}}$, revealing three distinct phases: a stable phase for $0<a_{\mathrm{fb}}\lesssim 4.0$, an unstable phase for $4.0\lesssim a_{\mathrm{fb}}\lesssim 8.0$, and an over-rotation phase for $8.0\lesssim a_{\mathrm{fb}}$. In the stable phase, $C_{\Sigma}$ shows an increasing trend with $a_{\mathrm{fb}}$, which indicates that the stronger feedback more effectively revives older information. This phase represents a well-balanced state where the feedback mechanism enhances memory retention capabilities without excessively diminishing the influence of the new input. In contrast, the unstable phase is characterized by a decline in $C_{\Sigma}$ as feedback intensifies, due to the system being predominantly governed by feedback rather than input. Such feedback dominance compromises the reservoir’s functionality in processing and preserving information effectively. The over-rotation phase is marked by a complete collapse of $C_{\Sigma}$ to zero, triggered by an excessively large $a_{\mathrm{fb}}$ leading to an over-rotation upon the feedback. Consequently, optimal memory performance is achieved with feedback strength just below the threshold for the onset of unstable behavior [36]. This point, in close proximity to the boundary between the stable and unstable phases, corresponds to the edge of chaos in this system.
@@ III.2 Measurement trajectories
However, this attractor-shaped dynamics collapses with further intensified feedback. Figures 4(c) and 4(d) illustrate typical dynamics in the unstable phase at $a_{\mathrm{fb}}=6.0$. Here, the dominant feedback causes the system to cycle through multiple stable states in the feature space, where the trajectory converges into a limit cycle. This undermines the influence of the input on the dynamics and, consequently, diminishes the system’s responsiveness to new inputs, leading to a decline in memory performance as $C_{\Sigma}\simeq 0$. Importantly, the formation of limit cycles depends on the specific realization of $U_{\mathrm{res}}$. For a particular $U_{\mathrm{res}}$ where the dynamics avoids entrapment in a limit cycle, attractor-like trajectories resembling those in Figs. 4(a) and 4(b) still emerge with a nonzero $C_{\Sigma}$ even at $a_{\mathrm{fb}}=6.0$. Nonetheless, as the feedback strength intensifies, the system exhibits an increased tendency to get caught in the limit cycle at a variety of $U_{\mathrm{res}}$ due to augmented instability. This results in a monotonic decrease in the capacity $C_{\Sigma}$ averaged over multiple instances of $U_{\mathrm{res}}$ in the unstable phase. In the over-rotation phase, the trajectory degrades into a random dynamics, with neither input nor feedback having a discernible influence [Figs. 4(e) and 4(f)]. These analysis underscore the significant influence of feedback connections on the dynamics of the quantum reservoir system, and thereby on the overall computational capabilities.
@@ IV.1 Dynamics of classical and quantum systems
Here, ${\sigma}_{i}^{x}$ and ${\sigma}_{i}^{z}$ denote the $x$ and $z$ Pauli matrices at the site $i$, while $h_{x}$ and $h_{z}$ signify the transverse and longitudinal magnetic fields, respectively. $J>0$ corresponds to the strength of the ferromagnetic nearest-neighbor Ising interaction, normalized to $J=1$ for our energy scale. Our focus lies in predicting the expectation value of the central spin’s $z$-component, $\langle\sigma^{z}_{3}(t)\rangle$, where the quantum system is initialized in the all up state and evolving according to the Heisenberg equation under the Hamiltonian in Eq. (8). In the absence of a longitudinal field ($h_{z}=0$), the spin system can be mapped to a free fermion system via the Jordan-Wigner transformation, leading to integrable dynamics [55]. Conversely, at $(h_{x},h_{z})=(1.05,-0.5)$, the system exhibits nonintegrable dynamics characterized by the chaotic spectral statistics [56]. The time-series $\{s_{k}\}$ is constructed by mapping the dynamics of the quantum spin from $t$ to $t+0.05$ to a single step from $k$ to $k+1$; visual representations of these inputs are provided in Fig. 5(a).
@@ IV.2 Predictive performance of the QRC
Figure 5: (a) Visualization of the time-series $\{s_{k}\}$ derived from classical MG dynamics, integrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$, and nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.05,-0.5)$. Each $\{s_{k}\}$ is depicted in the $s_{k}$-$s_{k+10}$ plane (top) and as a function of the time step $k$ (bottom). (b)-(e) Performance of the QRC and the ESN for the predictions of (b) cosine wave, (c) classical MG time-series, (d) integrable quantum spin dynamics, and (e) nonintegrable quantum spin dynamics. NMSE is averaged for $128$ random realizations for both the QRC and ESN. The QRC with $8$ qubits is constructed at $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{back}}=2.5$, and the ESN is equipped with $8$, $50$, and $100$ nodes. The performance of the ESN represents the lowest NMSE achieved through the optimization of network parameters.
@@
Figure 6: Measurement trajectories for a typical $U_{\mathrm{res}}$ in $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$, $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$, $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$, and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes. Each point represents $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ at cycle $k$, and each line connect $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ and $\left[\langle Z_{i}\rangle_{k+1},\langle Z_{i+1}\rangle_{k+1}\right]$. The input and feedback strength are set to $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{back}}=2.5$, respectively. The input sequence $\{s_{k}\}$ is derived from (a) cosine wave, (b) classical MG time-series, (c) integrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$, and (d) nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.05,-0.5)$.
@@
In Figs. 5(b)-5(e), we present the predictive performance of our QRC across four different time-series: a cosine wave $\{s_{k}\}=\{\cos(\omega k)\}$ with $\omega=\pi/25$, the classical MG time-series with $t_{d}=17$, and both the integrable and nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$ and $(h_{x},h_{z})=(1.05,-0.5)$, respectively. We employ the input and feedback weights of $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{fb}}=2.5$, ensuring that the system remains in the stable phase when subjected to random inputs (Figs. 3 and 4). For comparison, we present the performance of an echo state network (ESN), a typical classical reservoir computing framework, while varying the number of nodes and optimizing network parameters for each $\tau_{f}$; see Appendix A for further details.
@@
In predicting the cosine wave, the QRC exhibits remarkable performance, surpassing or comparable to the $100$-node ESN; $\mathrm{NMSE}\simeq 10^{-9}$ is achieved across all $\tau_{f}$ [Fig. 5(b)]. Moreover, it demonstrates pronounced predictive accuracy in forecasting both integrable and nonintegrable quantum spin dynamics, outperforming the $100$-node ESN up to $\tau_{f}\simeq 15$ [Figs. 5(d) and 5(e)]. (For $\tau_{f}>15$, both the QRC and the ESN perform poorly with $\mathrm{NMSE}>10^{-3}$, making comparisons in this regime less meaningful.) In contrast, for predictions of the classical chaotic MG dynamics, the effectiveness of our QRC diminishes, falling below the performance of both the $50$-node and $100$-node ESN for $\tau_{f}>3$, as illustrated in Fig. 5(c). These results highlight the strengths and limitations of our QRC in predicting diverse categories of time-series data: it shows a distinct advantage in predicting time-series originating from quantum spin dynamics, yet is less adept at predicting classical MG dynamics.
@@
Notably, our eight-qubit QRC consistently outperforms the ESN with an equivalent number of nodes and, for specific targets, even surpasses the $100$-node ESN. It is worth highlighting that our system necessitates the optimization of only $8+1$ parameters at the output part (including the bias term), which is a significant reduction compared to the $100+1$ optimization parameters in the $100$-node ESN. The remarkable performance of the QRC, despite having fewer computational nodes, can be attributed to the high-dimensionality of the quantum reservoir’s Hilbert space, widely recognized as a highly advantageous characteristic for effective reservoir computing [13, 19]. The reduction in computational nodes also contribute to the superior performance of the QRC at $\tau_{f}=0$; in such scenarios, where the objective is simply to reproduce the most recent input, excessive complexity can be detrimental.
@@
Figure 6 illustrates the measurement trajectories under the influence of each time-series signal $\{s_{k}\}$; for a typical realization of $U_{\mathrm{res}}$, $\left[\langle Z_{1}\rangle_{k},\langle Z_{2}\rangle_{k}\right]$, $\left[\langle Z_{3}\rangle_{k},\langle Z_{4}\rangle_{k}\right]$, $\left[\langle Z_{5}\rangle_{k},\langle Z_{6}\rangle_{k}\right]$, and $\left[\langle Z_{7}\rangle_{k},\langle Z_{8}\rangle_{k}\right]$ are depicted. When subjected to the cosine wave input, the dynamics of the quantum reservoir system exhibits periodic behavior, manifested as elliptical orbital trajectories in the feature space [Fig. 6(a)]. These trajectories possess a periodicity of $50$ cycles, which is consistent with the period of the original cosine wave input. This observation indicates that the internal dynamics of the quantum reservoir adequately reflects the provided input dynamics.
@@
In Fig. 6(b), we present the measurement outcomes influenced by the MG dynamics input. These trajectories, presumably reflecting the chaotic attractor observed in Fig. 5(a), form a complicated manifold that cannot be easily reduced to a simple two-dimensional representation. Conversely, under the input signals derived from integrable [Fig. 6(c)] and nonintegrable [Fig. 6(d)] quantum spin dynamics, the trajectories collapse into single lines in specific two-dimensional planes: the $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$, $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$, and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes in Fig. 6(c), and the $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ and $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$ planes in Fig. 6(d). Even when trajectories display broader dispersions, they predominantly manifest as superpositions of elliptical curves, each with varying radii and central locations [the $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ plane in Fig. 6(c), and the $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$ and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes in Fig. 6(d)]. This rotation-type reservoir dynamics is attributed to the characteristics of the provided signals, which originate from quantum spin dynamics that fundamentally involves spin rotations.

### v2

@@ III.1 Short-term memory capacity
Figure 3: (a) Reservoir performance for the short-term memory task $R^{2}_{d}$ plotted as a function of the delay $d$ with $a_{\mathrm{fb}}=2.0$ (blue), $a_{\mathrm{fb}}=4.0$ (green), $a_{\mathrm{fb}}=6.0$ (purple), and $a_{\mathrm{fb}}=8.0$ (yellow). (b) The dependency of the total capacity $C_{\Sigma}$ on the feedback strength $a_{\mathrm{fb}}$. Both $R^{2}_{d}$ and $C_{\Sigma}$ are averaged over $128$ samples with respect to $U_{\mathrm{res}}$.
Figure 3(a) exhibits the memory performance $R^{2}_{d}$ for varying feedback weights $a_{\mathrm{fb}}$ with the input weight fixed at $a_{\mathrm{in}}=0.001$. At the feedback strength of $a_{\mathrm{fb}}=2.0$, $R^{2}_{d}$ becomes close to $1$ for delays up to $d=4$, followed by a gradual decline for longer delays. Considering that the quantum state of the $k$-th cycle is determined only by $s_{k}$ and $\bm{z}_{k-1}$, this robust memory performance indicates that the input histories of $\{s_{k}\}$ are adequately accumulated within the measurement results $\bm{z}_{k-1}$. The quantum reservoir effectively extracts key features of the provided dataset through the read-out measurements, and these features are succeeded over different cycles via the feedback connections. In this manner, our feedback-driven QRC overcomes the destruction of the quantum state caused by the measurement back-action, thereby acquiring the fading-memory property.
@@
Figure 3(b) depicts the total capacity $C_{\Sigma}$ as a function of the feedback strength $a_{\mathrm{fb}}$, revealing three distinct phases: a stable phase for $0<a_{\mathrm{fb}}\lesssim 4.0$, an unstable phase for $4.0\lesssim a_{\mathrm{fb}}\lesssim 8.0$, and an over-rotation phase for $8.0\lesssim a_{\mathrm{fb}}$. In the stable phase, $C_{\Sigma}$ shows an increasing trend with $a_{\mathrm{fb}}$, which indicates that the stronger feedback more effectively revives older information. This phase represents a well-balanced state where the feedback mechanism enhances memory retention capabilities without excessively diminishing the influence of the new input. In contrast, the unstable phase is characterized by a decline in $C_{\Sigma}$ as feedback intensifies, due to the system being predominantly governed by feedback rather than input. Such feedback dominance compromises the reservoir’s functionality in processing and preserving information effectively. The over-rotation phase is marked by a complete collapse of $C_{\Sigma}$ to zero, triggered by an excessively large $a_{\mathrm{fb}}$ leading to the over-rotation upon the feedback. Consequently, optimal memory performance is achieved with feedback strength just below the threshold for the onset of unstable behavior [39]. This point, in close proximity to the boundary between the stable and unstable phases, corresponds to the edge of chaos in this system.
@@ III.2 Measurement trajectories
However, the attractor-shaped dynamics collapses with further intensified feedback. Figures 4(c) and 4(d) illustrate typical dynamics in the unstable phase at $a_{\mathrm{fb}}=6.0$. Here, the dominant feedback causes the system to cycle through multiple stable states in the feature space, where the trajectory converges into a limit cycle. This undermines the influence of the input on the dynamics and, consequently, diminishes the system’s responsiveness to new inputs, leading to a degradation in memory performance to $C_{\Sigma}\simeq 0$. Importantly, the formation of limit cycles depends on the specific realization of $U_{\mathrm{res}}$. For a particular $U_{\mathrm{res}}$ where the dynamics avoids entrapment in a limit cycle, attractor-like trajectories resembling those in Figs. 4(a) and 4(b) still emerge with a nonzero $C_{\Sigma}$ even at $a_{\mathrm{fb}}=6.0$. Nonetheless, as the feedback strength intensifies, the system exhibits an increased tendency to get caught in the limit cycle at a variety of $U_{\mathrm{res}}$ due to augmented instability. This results in a monotonic decrease in the capacity $C_{\Sigma}$ averaged over multiple instances of $U_{\mathrm{res}}$ in the unstable phase. In the over-rotation phase, the trajectory degrades into a random dynamics, with neither input nor feedback having a discernible influence [Figs. 4(e) and 4(f)].
As these findings indicate, the dynamics of the QRC, and consequently its overall computational capabilities, are determined by the balance between the input and feedback operations. It is important to note that, in the presence of statistical uncertainty in measurements, the performance is similarly governed by the interplay among the input, feedback, and statistical fluctuations. While such fluctuations can perturb the dynamics of the QRC to a certain degree, their effects can be mitigated by flexibly adjusting the input and feedback strengths. Therefore, the overall behavior of our QRC system remains largely unchanged even when measurement fluctuations are considered; see Appendix C for a more detailed analysis.
@@ IV.1 Dynamics of classical and quantum systems
Here, ${\sigma}_{i}^{x}$ and ${\sigma}_{i}^{z}$ denote the $x$ and $z$ Pauli matrices at the site $i$, while $h_{x}$ and $h_{z}$ signify the transverse and longitudinal magnetic fields, respectively. $J>0$ corresponds to the strength of the ferromagnetic nearest-neighbor Ising interaction, normalized to $J=1$ for our energy scale. Our focus lies in predicting the expectation value of the central spin’s $z$-component, $\langle\sigma^{z}_{3}(t)\rangle$, where the quantum system is initialized in the all up state and evolving under the Hamiltonian in Eq. (8). In the absence of a longitudinal field ($h_{z}=0$), the spin system can be mapped to a free fermion system via the Jordan-Wigner transformation, leading to integrable dynamics [57]. Conversely, at $(h_{x},h_{z})=(1.05,-0.5)$, the system exhibits nonintegrable dynamics characterized by the chaotic spectral statistics [58]. The time-series $\{s_{k}\}$ is constructed by mapping the dynamics of the quantum spin from $t$ to $t+0.05$ to a single step from $k$ to $k+1$; visual representations of these inputs are provided in Fig. 5(a).
@@ IV.2 Predictive performance of the QRC
Figure 5: (a) Visualization of the time-series input signal $\{s_{k}\}$ derived from classical MG dynamics, integrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$, and nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.05,-0.5)$. Each $\{s_{k}\}$ is depicted in the $s_{k}$-$s_{k+10}$ plane (top) and as a function of the time step $k$ (bottom). (b)-(e) Performance of the QRC and the ESN for the predictions of (b) cosine wave, (c) classical MG time-series, (d) integrable quantum spin dynamics, and (e) nonintegrable quantum spin dynamics. NMSE is averaged for $128$ random realizations for both the QRC and ESN. The QRC with $8$ qubits is constructed at $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{back}}=2.5$, and the ESN is equipped with $8$, $50$, and $100$ nodes. The performance of the ESN represents the lowest NMSE achieved through the optimization of network parameters.
@@
Figure 6: Measurement trajectories for a typical $U_{\mathrm{res}}$ in $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$, $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$, $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$, and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes. Each point represents $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ at cycle $k$, and each line connects $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ and $\left[\langle Z_{i}\rangle_{k+1},\langle Z_{i+1}\rangle_{k+1}\right]$. The input and feedback strength are set to $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{back}}=2.5$, respectively. The input sequence $\{s_{k}\}$ is derived from (a) cosine wave, (b) classical MG time-series, (c) integrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$, and (d) nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.05,-0.5)$.
@@
In Figs. 5(b)-5(e), we present the predictive performance of our QRC across four different time-series: a cosine wave $\{s_{k}\}=\{\cos(\omega k)\}$ with $\omega=\pi/25$, the classical MG time-series with $t_{d}=17$, and both the integrable and nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$ and $(h_{x},h_{z})=(1.05,-0.5)$, respectively. We employ the input and feedback weights of $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{fb}}=2.5$, ensuring that the system remains in the stable phase when subjected to random inputs (Figs. 3 and 4). For comparison, we present the performance of an echo state network (ESN), a typical classical reservoir computing framework. We optimize network parameters for each $\tau_{f}$ while changing the number of nodes; see Appendix D for further details.
@@
In predicting the cosine wave, the QRC exhibits remarkable performance, surpassing or comparable to the $100$-node ESN; $\mathrm{NMSE}\simeq 10^{-9}$ is achieved across all $\tau_{f}$ [Fig. 5(b)]. Moreover, it demonstrates pronounced predictive accuracy in forecasting both integrable and nonintegrable quantum spin dynamics, outperforming the $100$-node ESN up to $\tau_{f}\simeq 15$ [Figs. 5(d) and 5(e)]. (For $\tau_{f}>15$, both the QRC and the ESN perform poorly with $\mathrm{NMSE}>10^{-3}$, making comparisons in this regime less meaningful.) In contrast, for predictions of the classical chaotic MG dynamics, the effectiveness of our QRC diminishes, falling below the performance of both the $50$-node and $100$-node ESN for $\tau_{f}>3$ as illustrated in Fig. 5(c). These results highlight the strengths and limitations of our QRC in predicting diverse categories of time-series data: it shows a distinct advantage in predicting time-series originating from quantum spin dynamics, yet is less adept at predicting classical MG dynamics.
@@
Notably, our eight-qubit QRC consistently outperforms the ESN with an equivalent number of nodes and, for specific targets, even surpasses the $100$-node ESN. Corresponding to the number of qubits, our system necessitates the optimization of only $8+1$ parameters at the output part (including the bias term), which is a significant reduction compared to the $100+1$ optimization parameters in the $100$-node ESN. The remarkable performance of the QRC despite having fewer computational nodes can be attributed to the high-dimensionality of the quantum reservoir’s Hilbert space, widely recognized as a highly advantageous characteristic for effective reservoir computing [14, 20]. The reduction in computational nodes also contribute to the superior performance of the QRC at $\tau_{f}=0$; in such scenarios, where the objective is simply to reproduce the most recent input, excessive complexity can be detrimental.
In Appendix E, we present the predictive performance of the QRC across various system sizes, utilizing the input signal derived from the nonintegrable quantum spin dynamics. Our proposed protocol can be seamlessly extended to different system sizes with analogous circuit architectures as depicted in Fig. 2(a). Notably, our QRC demonstrates superior performance compared to the $100$-node ESN even when utilizing only $N=6$ qubits, further underscoring the effectiveness of our approach in processing this quantum-origin input signal. In addition, the observed monotonic improvement in performance with increasing system size provides compelling evidence for the potential of further enhancements in larger systems.
@@
Figure 6 illustrates the measurement trajectories under the influence of each time-series signal $\{s_{k}\}$; for a typical realization of $U_{\mathrm{res}}$, $\left[\langle Z_{1}\rangle_{k},\langle Z_{2}\rangle_{k}\right]$, $\left[\langle Z_{3}\rangle_{k},\langle Z_{4}\rangle_{k}\right]$, $\left[\langle Z_{5}\rangle_{k},\langle Z_{6}\rangle_{k}\right]$, and $\left[\langle Z_{7}\rangle_{k},\langle Z_{8}\rangle_{k}\right]$ are depicted. When subjected to the cosine wave input, the dynamics of the quantum reservoir system exhibits periodic behavior, manifested as elliptical orbital trajectories in the feature space [Fig. 6(a)]. These trajectories possess a periodicity of $50$ cycles, which is consistent with the period of the original cosine wave input. This observation indicates that the internal dynamics of the quantum reservoir adequately reflects the characteristics of the provided input dynamics.
@@
In Fig. 6(b), we present the measurement outcomes influenced by the MG dynamics input. These trajectories, presumably reflecting the chaotic attractor observed in Fig. 5(a), form a complicated manifold that cannot be easily reduced to a simple two-dimensional representation. Conversely, under the input signals derived from integrable [Fig. 6(c)] and nonintegrable [Fig. 6(d)] quantum spin dynamics, the trajectories collapse into single lines in specific two-dimensional planes: the $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$, $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$, and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes in Fig. 6(c), and the $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ and $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$ planes in Fig. 6(d). Even when trajectories display broader dispersions, they predominantly manifest as superpositions of elliptical curves, each with varying radii and central locations [the $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ plane in Fig. 6(c), and the $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$ and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes in Fig. 6(d)]. This rotation-type reservoir dynamics is attributed to the properties of the provided signals, which originate from quantum spin dynamics that fundamentally involves spin rotations.